OpenStudy (anonymous):

(5+5i)(8(cos((3pi)/7)+(i)sin((3pi)/7))

OpenStudy (anonymous):

@ganeshie8 @nikato @jim_thompson5910

random231 (random231):

http://mathworld.wolfram.com/ComplexNumber.html

OpenStudy (anonymous):

@random231 lol that didn't help...like at all...

ganeshie8 (ganeshie8):

\[\large (5+5i)(8(\cos((3\pi)/7)+i\sin((3\pi)/7)) = (5+5i)e^{i\frac{3\pi}{7}}\]

random231 (random231):